# µnit Scaling: Simple and Scalable FP8 LLM Training

### v2

@@ Abstract
Large Language Model training with 8-bit floating point (FP8) formats promises significant efficiency improvements, but reduced numerical precision makes training challenging. It is currently possible to train in FP8 only if one is willing to tune various hyperparameters, reduce model scale, or accept the overhead of computing dynamic scale factors.
We demonstrate simple, scalable FP8 training that requires no dynamic scaling factors or special hyperparameters, even at large model sizes. Our method, µnit Scaling (µS), also enables simple hyperparameter transfer across model widths, matched numerics across training and inference, and other desirable properties. µnit Scaling is straightforward to implement, consisting of a set of minimal interventions based on a first-principles analysis of common transformer operations.
We validate our method by training models from 1B to 13B parameters, performing all hidden linear layer computations in FP8. We achieve quality equal to higher precision baselines while also training up to 33% faster.
@@ 1 Introduction
Because LLM training is computationally expensive, low-precision training provides large compute savings. Modern LLMs are typically trained in mixed-precision bfloat16 (BF16), where most computation occurs in BF16, but some components requiring higher precision (such as accumulators and master weights) use FP32 Micikevicius et al. (2018). Thanks to increased hardware support for FP8 formats, mixed precision training using FP8 computation promises even greater training efficiency Micikevicius et al. (2022). However, the reduced range and resolution of FP8 make LLM training challenging. In this work, we demonstrate a simple, scalable FP8 training method with straightforward hyperparameter transfer on large LLMs, called “µnit Scaling” (µS).
Our µnit Scaling method builds on Unit Scaling Blake et al. (2023), which aims to maintain unit variance in weights, activations, and gradients. To ensure this, it scales neural network operations with static constants and initializes network parameters to have unit variance. If all tensors used in training can maintain unit variance, they are representable with sufficient range and resolution by low-precision formats like FP16 and FP8. However, preserving high-quality tensor representations in low-precision formats is challenging for large models.
@@
Straightforward hyperparameter transfer Tuning hyperparameters for large LLMs is expensive. A promising way to reduce this cost is to tune the hyperparameters for smaller LLMs and “transfer” them to large ones, either by using them directly or by applying a model-size-based formula as explored in µ-Parametrization (µP)Yang et al. (2021; 2024a; 2024b). However, applying hyperparameter transfer techniques in practice to low-precision training can be challenging; frequent divergences due to numerical issues may require training in higher precisions like FP32 Yang et al. (2021). To address this, Blake et al. (2024) introduced u-µP, which combines Unit Scaling Blake et al. (2023) and µP to enable hyperparameter transfer in low precision. Unfortunately, compared to conventional BF16 mixed precision training (henceforth termed “standard parametrized” (SP) models), both µP and u-µP have many more hyperparameters to sweep over as shown in Table 3, diminishing realized compute savings and increasing complexity. Specific implementation intricacies, such as zero-initialized queries in µP or LR scaling for embeddings by fan-out in u-µP, make these schemes harder to use in practice than SP. In contrast, our µnit Scaling (µS) scheme combines µP and Unit Scaling in a greatly simplified way, making it easier to use and more cost-effective. We demonstrate hyperparameter transfer of learning rate ($\eta$) and weight decay ($\lambda$) to models of up to 20x larger widths.
@@
No Dynamic Scaling With dynamic scaling, one calculates per-tensor scaling factors for each weight, activation, and gradient tensor in training. These scales shift BF16 tensors into the representable ranges of FP8 formats in each forward and backward pass. Typically, one also decouples the forward and backward formats, using e4m3 for weights and activations and e5m2 for gradients Sun et al. (2019). NVIDIA’s TransformerEngine is a notable example of an FP8 training library that uses dynamic scaling NVIDIA (2023). Calculating scaling factors dynamically adds training and inference overhead and complicates large-scale distributed training and checkpointing.
@@
Apply to All Linear Layers Existing work on applying Unit Scaling at larger scales requires certain “critical matmuls” (attention out projection, FFN down projection) to stay in BF16 Blake et al. (2024). Assuming a transformer model with conventional multiheaded attention and an MLP with an expansion ratio of 4, this means 41.7% of all hidden linear layer FLOPs are not in FP8. In contrast, µS ensures that, regardless of scale, all hidden layers use FP8.
@@
Match Inference-Time Quantization For efficient inference, LLMs are often quantized to FP8 or INT8 for faster computation and reduced memory footprints Khudia et al. (2021); Dettmers et al. (2022). Since training typically occurs in higher bitwidths (e.g., BF16), a mismatch in precisions at training time and inference time means that some level of quantization error is unavoidable, degrading model quality. Training with µS avoids this mismatch—since the LLM has already been trained in FP8, it is immediately ready for inference in FP8 for both weights and activations (W8A8).
@@ 2 Methods
Table 1: Components of the µS training scheme. µS makes the following modifications to standard decoder-only transformer practices.
@@ 2.1 Self-attention Numerics
$\begin{split}\mu_{\mathbf{n}}=\mathrm{e}^{1/2},\quad\sigma^{2}_{\mathbf{n}}=\mathrm{e}(\mathrm{e}-1)\\ \mu_{\mathbf{d}}=k\mathrm{e}^{1/2},\quad\sigma^{2}_{\mathbf{d}}=k\mathrm{e}(\mathrm{e}-1)\\ \mathrm{Cov}[\mathbf{n},\mathbf{d}]=\sigma^{2}_{\mathbf{n}}=\mathrm{e}(\mathrm{e}-1)\end{split}$ (2)
@@
$\begin{split}\sigma^{2}_{\mathbf{s}}=\mathrm{Var}\left[\frac{\mathbf{n}}{\mathbf{d}}\right]&\approx\frac{\sigma^{2}_{\mathbf{n}}}{\mu^{2}_{\mathbf{d}}}+\frac{\mu^{2}_{\mathbf{n}}\sigma^{2}_{\mathbf{d}}}{\mu^{4}_{\mathbf{d}}}-2\frac{\mu_{\mathbf{n}}\mathrm{Cov}[\mathbf{n},\mathbf{d}]}{\mu^{3}_{\mathbf{d}}}\\ &=\frac{\mathrm{e}-1}{k^{2}}-\frac{\mathrm{e}-1}{k^{3}}\end{split}$ (4)
@@
$\sigma^{2}_{\mathbf{a}}=\sum_{i=1}^{k}\sigma^{2}_{\mathbf{s}}\sigma^{2}_{\mathbf{V}}+\sigma^{2}_{\mathbf{s}}\mu^{2}_{\mathbf{V}}+\sigma^{2}_{\mathbf{V}}\mu^{2}_{\mathbf{s}}=\frac{\mathrm{e}}{k}-\frac{\mathrm{e}-1}{k^{2}}$ (6)
@@
That is, by taking the square root of attention scores, attention can be made variance-preserving for independent value tokens. This modification, which we term “Square-Root Softmax attention”, is shown in Eq. 9. Square-Root Softmax attention is also easily implemented via modern attention kernels like Flex-Attention Dong et al. (2024).
@@
Figure 3: Value tokens in text are highly correlated. Comparison of cosine similarity between observed value tokens in a text data distribution versus value tokens $\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,1)$. Repeated tokens in the value matrix, an unavoidable result of token frequency in real text data, lead to higher-than-random $\sigma$ as sequence position increases (c.f. Fig. 2).
Figure 4: Res-Post-LayerNorm. (a) Pre-LayerNorm transformer architecture versus Res-Post-LayerNorm architecture. Res-Post-LayerNorm moves the LayerNorm operation from the start of each residual branch to the end Liu et al. (2022). This ensures consistent variance across tokens when added to the residual stream. In contrast, Pre-LayerNorm networks permit unnormalized representations with inconsistent variance to be added to the residual stream, as shown with self-attention outputs in Fig. 2. (b) Convergence test loss curves with 100-layer models show that µS with Res-Post-LayerNorm achieves nearly identical convergence versus SP with Pre-LayerNorm.
@@ 2.2 Residual Modification Schemes
Every skip connection in a neural network adds another tensor into the residual stream. Summing all these tensors tends to increase the variance of the residual stream deeper in the network. To instead make residual connections variance-preserving, Blake et al. (2023) proposed to replace simple summation with weighted summation, where the weights $a$ and $b$ of the skip connection and residual branch satisfy $a^{2}+b^{2}=1$. They proposed two methods for setting these coefficients: fixed and running-mean, which are shown in Eq. 10 and Eq. 11 respectively. The former uses a constant coefficient $\tau$ while the latter uses coefficients which are a function of the layer index $l$.
@@
As shown in Fig. 5, we found that the fixed scheme provides better convergence than the running-mean scheme. All µS models we train therefore use the fixed scheme. We set the coefficient $\tau$ based on the depth using the results in Appendix A.2.
Figure 5: Residual Modification schemes affect µnit Scaled model convergence. The fixed residual modification (green, Eq. 10) achieves better training convergence for deep transformers than the running-mean residual modification (blue, Eq. 11). The fixed residual coefficient for this model is $\tau=0.1$.
@@ 2.3 Hyperparameter Transfer with µnit Scaling
Zero-shot hyperparameter transfer allows hyperparameters to be tuned on a small proxy network, then directly used on much larger networks without any further tuning Yang et al. (2021). The width of the small proxy network is typically referred to as the “base width”, or $d_{\text{base}}$. Because it eliminates the need to sweep hyperparameters at a large scale, such hyperparameter transfer yields massive compute savings.
Hyperparameter transfer with µnit Scaling follows from neural network equivalencies set forth in Yang et al. (2021, Appendix J.2.1), reproduced below for convenience. As detailed in Blake et al. (2024), Equations 1213, and 14 define the hidden layer in a model undergoing training. All hidden layers are initialized with weights $\mathbf{W}_{0}$ drawn from a normal distribution with variance $b^{2}$, use a learning rate of $c$, and have an output multiplier $a$. $\mathbf{X}$ and $\mathbf{Y}$ denote input and output activation matrices respectively; $t$ is the training time step; and $\mathbf{\Phi}_{t}(\nabla\mathcal{L}_{0},\ldots,\nabla\mathcal{L}_{t})$ denotes the weight update for time step $t$ using prior loss gradients.
@@
Adam-like $\lambda$ 1 1 $1/\sqrt{\text{fan\_in}}$
@@ 3.1 Successful Hyperparameter Transfer
Setup: To evaluate hyperparameter transfer, we first train four-layer decoder-only LLMs with widths of 256 through 8192 using Standard Parametrization (SP) and µnit Scaling (µS). We begin with these these small models since doing so allows us to collect ground truth optimal hyperparameters. All models use multi-headed attention Vaswani et al. (2017) and were trained for 10,000 training steps with a global batch size of 64 and sequence length of 1024 (i.e., 655M total tokens). SP models use Pre-LayerNorm placement and are trained in both BF16 and FP8 (using TransformerEngine). µS models were trained in both BF16 and FP8 and use Res-Post-LayerNorm placement (Fig. 4). µS used base models of width 256. For all models described in this and subsequent sections, we used the Lion optimizer Chen et al. (2023) with fully decoupled weight decay and a cosine learning rate schedule decaying to 10% of the maximum learning rate. For details on why Lion is an Adam-like optimizer for hyperparameter transfer, please refer to Appendix A.3. All models were trained on Nvidia H100 GPUs using the Databricks MosaicML LLMFoundry MosaicML (2022a), Composer MosaicML (2021), and Streaming MosaicML (2022b) libraries.
Hyperparameters: We evaluate hyperparameter transfer over learning rate ($\eta$) and weight decay ($\lambda$). While µP Yang et al. (2021) does not give a theoretical basis for $\lambda$ transfer over width, we evaluate its transfer empirically because of its practical importance. We use fully decoupled weight decay, motivated by findings from Wortsman et al. (2024), Lingle (2024), and Wang & Aitchison (2024). To elucidate how $\lambda$ scales with model width, we jointly sweep over both $\eta$ and $\lambda$ with powers of 2. Based on the relationship between the residual coefficient $\tau$ and depth in Appendix A.2, the residual coefficient $\tau$ is 0.4 for these four layer models.
@@ 3.2 FP8 Training at Scale
Model Param. Tokens TPR Steps Batch Sz. Seq. Len. Width Depth # Heads $\tau$
@@
Setup: We train 1B, 3B, 7B, and 13B parameter LLMs on approximately compute-optimal token budgets ($\sim$20x token to parameter ratio) using SP and µS, and in both BF16 and FP8, resulting in 4 individual models for each model size. The training configurations are detailed in Table 4. Based on the previous sections’ hyperparameter transfer results (Fig. 6), we sweep $\eta$ and $\lambda$ on small models with a base width of $d_{\text{base}}=256$, then transfer optimal hyperparameters to large models with width $d_{\text{new}}$, as shown below.
@@
Evaluation: We use the Databricks Model Gauntlet to evaluate the quality of all models on specific tasks Dohmann (2023); Barton (2024). These results are shown in Table 5.
@@
We also note that at the 13B scale, we attempted to remedy the divergence of the SP FP8 model by using multiple different values of $\lambda$, but this did not mitigate the frequent loss spikes and eventual divergence. µS models, by contrast, train stably.

### v3

@@ Abstract
Large language model training with 8-bit floating point (FP8) formats promises significant efficiency improvements, but reduced numerical precision makes training challenging. It is currently possible to train in FP8 only if one is willing to tune various hyperparameters, reduce model scale, or accept the overhead of computing dynamic scale factors. We demonstrate simple, scalable FP8 training that requires no dynamic scaling factors or special hyperparameters, even at large model sizes. Our method, µnit Scaling (µS), also enables simple hyperparameter transfer across model widths, matched numerics across training and inference, and other desirable properties. µnit Scaling is straightforward to implement, consisting of a set of minimal interventions based on a first-principles analysis of transformer operations. We validate our method by training models with parameters ranging from 1B to 13B, performing all hidden linear layer computations in FP8. We achieve quality equal to higher-precision baselines while also training up to 33% faster.
LLM, FP8, Transformer, Model Training, Attention
@@ 1 Introduction
Because LLM training is computationally expensive, low-precision training provides large compute savings. Modern LLMs are typically trained in mixed-precision bfloat16 (BF16), where most computation occurs in BF16, but some components requiring higher precision (such as accumulators and master weights) use FP32 (Micikevicius et al., 2018). Thanks to increased hardware support for FP8 formats, mixed precision training using FP8 computation promises even greater training efficiency (Micikevicius et al., 2022). However, the reduced range and resolution of FP8 make LLM training challenging. In this work, we demonstrate a simple, scalable FP8 training method with straightforward hyperparameter transfer on large LLMs, called “µnit Scaling” (µS).
Our µnit Scaling method builds on Unit Scaling (Blake et al., 2023), which aims to maintain unit variance in weights, activations, and gradients. To ensure this, it scales neural network operations with static constants and initializes network parameters to have unit variance. If all tensors used in training can maintain unit variance, they are representable with sufficient range and resolution by low-precision formats like FP16 and FP8. However, preserving high-quality tensor representations in low-precision formats is challenging for large models.
@@
Straightforward hyperparameter transfer Tuning hyperparameters for large LLMs is expensive. A promising way to reduce this cost is to tune the hyperparameters for smaller LLMs and “transfer” them to large ones, either by using them directly or by applying a model-size-based formula as explored in µ-Parametrization (µP)(Yang et al., 2021, 2023, 2024). However, applying hyperparameter transfer techniques in practice to low-precision training can be challenging; frequent divergences due to numerical issues may require training in higher precisions like FP32 (Yang et al., 2021). To address this, Blake et al. (2024) introduced u-µP, which combines Unit Scaling (Blake et al., 2023) and µP to enable hyperparameter transfer in low precision. Unfortunately, compared to conventional BF16 mixed precision training (henceforth termed “standard parametrized” (SP) models), both µP and u-µP have many more hyperparameters to sweep over (see Table 3), diminishing realized compute savings and increasing complexity. Specific implementation intricacies, such as zero-initialized queries in µP or LR scaling for embeddings by fan-out in u-µP, make these schemes harder to use in practice than SP. In contrast, our µnit Scaling (µS) scheme combines µP and Unit Scaling in a greatly simplified way, making it easier to use and more cost-effective. We demonstrate hyperparameter transfer of learning rate ($\eta$) and weight decay ($\lambda$) to models of up to 20x larger widths.
@@
No Dynamic Scaling With dynamic scaling, one calculates per-tensor scaling factors for each weight, activation, and gradient tensor in training. These scales shift BF16 tensors into the representable ranges of FP8 formats in each forward and backward pass. Typically, one also decouples the forward and backward formats, using e4m3 for weights and activations and e5m2 for gradients (Sun et al., 2019). NVIDIA’s TransformerEngine is a notable example of an FP8 training library that uses dynamic scaling (NVIDIA, 2023). Calculating scaling factors dynamically adds training and inference overhead and complicates large-scale distributed training and checkpointing.
@@
Apply to All Linear Layers Existing work on applying Unit Scaling at larger scales requires certain “critical matmuls” (attention out projection, FFN down projection) to stay in BF16 (Blake et al., 2024). Assuming a transformer model with conventional multiheaded attention and an MLP with an expansion ratio of 4, this means 41.7% of all hidden linear layer FLOPs are not in FP8. In contrast, µS ensures that, regardless of scale, all hidden layers use FP8.
@@
Match Inference-Time Quantization For efficient inference, LLMs are often quantized to FP8 or INT8 for faster computation and reduced memory footprints (Khudia et al., 2021; Dettmers et al., 2022). Since training typically occurs in higher bitwidths (e.g., BF16), a mismatch in precisions at training time and inference time means that some level of quantization error is unavoidable, degrading model quality. Training with µS avoids this mismatch—since the LLM has already been trained in FP8, it is immediately ready for inference in FP8 for both weights and activations (W8A8).
@@ 2 Methods
Table 1: Components of the µS training scheme. µS makes the following modifications to standard decoder-only transformer training practices. A deeper explanation of these modifications is provided in Appendix A.1.
@@ 2.1 Self-attention Numerics
$\begin{split}\mu_{\mathbf{n}}=e^{1/2},\quad\sigma^{2}_{\mathbf{n}}=e(e-1)\\ \mu_{\mathbf{d}}=ke^{1/2},\quad\sigma^{2}_{\mathbf{d}}=ke(e-1)\\ \mathrm{Cov}[\mathbf{n},\mathbf{d}]=\sigma^{2}_{\mathbf{n}}=e(e-1)\end{split}$ (2)
@@
$\begin{split}\sigma^{2}_{\mathbf{s}}=\mathrm{Var}\left[\frac{\mathbf{n}}{\mathbf{d}}\right]&\approx\frac{\sigma^{2}_{\mathbf{n}}}{\mu^{2}_{\mathbf{d}}}+\frac{\mu^{2}_{\mathbf{n}}\sigma^{2}_{\mathbf{d}}}{\mu^{4}_{\mathbf{d}}}-2\frac{\mu_{\mathbf{n}}\mathrm{Cov}[\mathbf{n},\mathbf{d}]}{\mu^{3}_{\mathbf{d}}}\\ &=\frac{e-1}{k^{2}}-\frac{e-1}{k^{3}}\end{split}$ (4)
@@
$\sigma^{2}_{\mathbf{a}}=\sum_{i=1}^{k}\sigma^{2}_{\mathbf{s}}\sigma^{2}_{\mathbf{V}}+\sigma^{2}_{\mathbf{s}}\mu^{2}_{\mathbf{V}}+\sigma^{2}_{\mathbf{V}}\mu^{2}_{\mathbf{s}}=\frac{e}{k}-\frac{e-1}{k^{2}}$ (6)
@@
That is, by taking the square root of attention scores, attention can be made variance-preserving for independent value tokens. This modification, which we term “Square-Root Softmax attention”, is shown in Eq. 9. Square-Root Softmax attention is also easily implemented via modern attention kernels like Flex-Attention (Dong et al., 2024).
@@
Figure 3: Value tokens in text are highly correlated. Comparison of cosine similarity between observed value tokens in a text data distribution versus value tokens $\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,1)$. Repeated tokens in the value matrix, an unavoidable result of token frequency in real text data, lead to higher-than-random $\sigma$ as sequence position increases (cf. Fig. 2).
Figure 4: Res-Post-LayerNorm. (a) Pre-LayerNorm transformer architecture versus Res-Post-LayerNorm architecture. Res-Post-LayerNorm moves the LayerNorm operation from the start of each residual branch to the end (Liu et al., 2022). This ensures consistent variance across tokens when added to the residual stream. In contrast, Pre-LayerNorm networks permit unnormalized representations with inconsistent variance to be added to the residual stream, as shown with self-attention outputs in Fig. 2. (b) Convergence test loss curves with 100-layer models show that µS with Res-Post-LayerNorm achieves nearly identical convergence versus SP with Pre-LayerNorm. (c) Additional convergence tests with 100-layer models show that Res-Post-LayerNorm achieves better convergence over Pre-LayerNorm with µS.
@@ 2.2 Residual Modification Schemes
Every skip connection in a neural network adds another tensor to the residual stream. Summing all these tensors tends to increase the variance of the residual stream deeper in the network. To make residual connections variance-preserving instead, Blake et al. (2023) proposed replacing simple summation with weighted summation, where the weights $a$ and $b$ of the skip connection and residual branch satisfy $a^{2}+b^{2}=1$. They proposed two methods for setting these coefficients: fixed and running-mean, which are shown in Eq. 11 and Eq. 12, respectively. The former uses a constant coefficient $\tau$, while the latter uses coefficients that are a function of the layer index $l$. The standard residual layer modification is shown in Eq. 10.
$\text{standard}:x_{l+1}=x_{l}+f(x_{l})$ (10)
@@
As shown in Fig. 5, we found that using either modification is better than the standard approach, with the fixed scheme providing better convergence than the running-mean scheme. All µS models we train therefore use the fixed scheme. We set the coefficient $\tau$ based on the depth using the results in Appendix A.3.
Figure 5: Residual modification schemes affect µnit Scaled model convergence. The fixed residual modification (green, Eq. 11) achieves better training convergence for deep transformers than the running-mean residual modification (blue, Eq. 12). The fixed residual coefficient for this model is $\tau=0.1$. Both of these settings outperform the standard residual layer modification (red, Eq. 10).
@@ 2.3 Hyperparameter Transfer with µnit Scaling
Zero-shot hyperparameter transfer allows hyperparameters to be tuned on a small proxy network, then directly used on much larger networks without any further tuning (Yang et al., 2021). The width of the small proxy network is typically referred to as the “base width”, or $d_{\text{base}}$. Because it eliminates the need to sweep hyperparameters at a large scale, such hyperparameter transfer yields massive compute savings.
Hyperparameter transfer with µnit Scaling follows from neural network equivalencies set forth in Yang et al. (2021, Appendix J.2.1), reproduced below for convenience. As detailed in Blake et al. (2024), Equations 1314, and 15 define the hidden layer in a model undergoing training. All hidden layers are initialized with weights $\mathbf{W}_{0}$ drawn from a normal distribution with variance $b^{2}$, use a learning rate of $c$, and have an output multiplier $a$. $\mathbf{X}$ and $\mathbf{Y}$ denote input and output activation matrices respectively; $t$ is the training time step; and $\mathbf{\Phi}_{t}(\nabla\mathcal{L}_{0},\ldots,\nabla\mathcal{L}_{t})$ denotes the weight update for time step $t$ using prior loss gradients.
@@
Adam-like LR 1 1 $1/\sqrt{\text{fan\_in}}$
@@ 3.1 Successful Hyperparameter Transfer
Setup: To evaluate hyperparameter transfer, we first train four-layer decoder-only LLMs with widths of 256 through 8192 using Standard Parametrization (SP) and µnit Scaling (µS). We begin with these small models since doing so allows us to collect ground truth optimal hyperparameters. All models use multi-headed attention (Vaswani et al., 2017) and were trained for 10,000 training steps with a global batch size of 64 and sequence length of 1024 (i.e., 655M total tokens). SP models use Pre-LayerNorm placement and are trained in both BF16 and FP8 (using TransformerEngine). µS models were trained in both BF16 and FP8 and use Res-Post-LayerNorm placement (Fig. 4). µS used base models of width 256. For all models described in this and subsequent sections, we used the Lion optimizer (Chen et al., 2023) with fully decoupled weight decay and a cosine learning rate schedule decaying to 10% of the maximum learning rate. For details on why Lion is an Adam-like optimizer for hyperparameter transfer, please refer to Appendix A.4. All models were trained on Nvidia H100 GPUs using the Databricks MosaicML LLMFoundry (MosaicML, 2022a), Composer (MosaicML, 2021), and Streaming (MosaicML, 2022b) libraries.
Hyperparameters: We evaluate hyperparameter transfer over learning rate ($\eta$) and weight decay ($\lambda$). While µP Yang et al. (2021) does not give a theoretical basis for $\lambda$ transfer over width, we evaluate its transfer empirically because of its practical importance. Prior work by Lingle (2024) has shown that µP does not admit transfer of $\lambda$ with AdamW. However, Wang & Aitchison (2024) found that optimal $\lambda$ should scale with model size. To elucidate how $\lambda$ scales with model width, we jointly sweep over both $\eta$ and $\lambda$. We use fully decoupled weight decay, motivated by findings from Wortsman et al. (2024) that doing so results in more stable training. $\eta$ and $\lambda$ are swept over powers of 2. Based on the relationship between the residual coefficient $\tau$ and depth in Appendix A.3, the residual coefficient $\tau$ is 0.4 for these four-layer models.
@@ 3.2 FP8 Training at Scale
Model Params Tokens TPR Steps Batch Sz. Seq. Len. Width Depth # Heads $\tau$
@@
Setup: We train 1B, 3B, 7B, and 13B parameter LLMs on approximately compute-optimal token budgets ($\sim$20x token-to-parameter ratio) using SP and µS, and in both BF16 and FP8, resulting in 4 individual models for each model size. The training configurations are detailed in Table 4. Based on the previous sections’ hyperparameter transfer results (Fig. 6), we sweep $\eta$ and $\lambda$ on small models with a base width of $d_{\text{base}}=256$, then transfer optimal hyperparameters to large models with width $d_{\text{new}}$, as shown below.
@@
Evaluation: We use the Databricks Model Gauntlet to evaluate the quality of all models on specific tasks (Dohmann, 2023; Barton, 2024). These results are shown in Table 5.
@@
We also note that at the 13B scale, we attempted to remedy the divergence of the SP FP8 model by using multiple different values of $\lambda$, but this did not mitigate the frequent loss spikes and eventual divergence. µS models, by contrast, train stably. We also show the instability in training with Unit Scaling (US) at larger scales in Appendix A.5, motivating runs only with SP and µS for our final results.